OpenStudy (anonymous):

Use an Addition or Subtraction to Formula to write the expression as a trigonometric function of one number, and then find its exact value.

OpenStudy (anonymous):

\[\cos (13\pi/15)\cos (-\pi/5)-\sin (13\pi/15)\sin (-\pi/5)\]